# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

### v1

@@ 3 Analog Pipeline Model Training
Model parallelism without pipeline. Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data, as previously mentioned) for iteration $k$. Vanilla model parallelism processes all micro-batches one by one. In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$, while stage $m$ achieves the backward error $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage in the backward pass. The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stored when the gradient for the $b$-th micro-batch is completed. After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method. Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}$, which iterates over $b\in[B_{\text{mini}}]$ by
@@
Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$. However, for the AIMC accelerators, only rank-update is permitted so that each gradient is accumulated to the weight $W^{(m)}_{k}$ one by one, rendering the analysis more difficult.
@@
Synchronous pipeline. Existing work has already attempted to design architecture that supports the pipeline training [24, 25]. They faithfully implement the iteration (7) and overlap parts of the computation. Since they compute all gradients in a batch on the same $W_{k}$, their algorithms can be called synchronous pipeline. In the vanilla non-pipeline version, the stage $m$ is idle after its forward computation of $b$-th micro-batch until the backward error comes back from stage $m+1$. In the synchronous pipeline training, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any; see the middle of Figure 2. Similar to the vanilla non-pipeline version, the weights will be updated after processing all micro-batches. To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch (fill and drain periods), hence rendering some devices idle.
Asynchronous pipeline. To increase the usage of all devices, a more progressive strategy, asynchronous pipeline, is proposed. Instead of waiting for the gradient of one mini-batch, the asynchronous pipeline training method performs gradient-based update algorithms immediately after one micro-batch is completed. Except for short periods of fill and drain at the beginning and the end, all devices are active during the training, maximizing the training efficiency; see the bottom of Figure 2.
@@
Despite its perfect resource usage, asynchronous pipeline parallelism suffers from the stale signal issue. Notice that in the asynchronous pipeline setting, the forward signal $x^{(m)}$ and backward signals $\delta^{(m)}$ are computed based on different stale weights. For example, after outputting forward signal $x^{(0)}$ for the $k$-th micro-batch, the first stage ($m=1$) receives the backward signal $\delta^{(0)}$ after $2(M-1)$ clock cycles, wherein $M-1$ gradient updates happen. Consequently, the gradient is stale for the current weights. Furthermore, notice both forward (2) and backward (3) passes involve the matrix $W^{(m)}$ while the $W^{(m)}$ is changed between the forward and backward passes. To depict the process more precisely, we introduce different symbols, which have tildes on top of them to indicate the staleness. Since the mini-batch is no longer needed in asynchronous pipeline, the subscript $b$ for the micro-batch index can be omitted, and we denote the micro-batch for $k$-th update as $\xi_{k}$. To make the relationship clearer, the computation dependency is plotted in Figure 3.
Since there are $M-m$ gradient updates between the forward and backward passes in stage $m$, the weight used for the forward pass has a delay of $M-m$, that is
${\tilde{x}}^{(m+1)}_{k}=$ $\ g^{(m)}({\tilde{z}}^{(m)}_{k}),\quad{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},\quad{\tilde{x}}^{(0)}_{k}=x_{k}.$ (8)
In contrast, the weights are updated following the backward pass, which means the weight used for backward is the latest one
$\tilde{\delta}^{(m)}_{k}=$ $\ \tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k}),$ (9)
@@
$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
@@
In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is no longer a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partial delay gradient, making the analysis more challenging.
Remark 1 (Digital v.s. analog asynchronous pipeline).
In digital hardware, one can stash the stale weight $W^{(m)}_{k-(M-m)}$ and read it out again to compute the backward signal in (9) [26, 21]. Weight stashing ensures the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is the gradient with respect to $W^{(m)}_{k-(M-m)}$, and hence provides a better convergence guarantee. However, copying is expensive in the analog domain, as we discussed before. Consequently, the asynchronous pipeline suffers from partial delay gradient issues.
Computation density. Different pipeline strategies can process varying amounts of data in the same amount of time. In an ideal scenario, a micro-batch requires $2M$ clock cycles for forward and backward passes, distributed equally across $M$ devices, taking 2 clock cycles per micro-batch. To measure how closely an algorithm approaches this ideal, we introduce computation density. Let $N_{\text{c}}(t)$ be the number of data points processed in $t$ clock cycles. Assuming infinite runtime, we define computation density as the limit ratio of ideal processing time to real time: $\lim_{t\to\infty}\frac{2N_{\text{c}}(t)}{t}$, where we introduce the factor $2$ since each data involves forward and backward passes. For vanilla model parallelism without pipelining, it takes $2M$ cycles per micro-batch, yielding a computation density of $1/M$. In synchronous pipelining, it takes $2(M+B-1)$ clock cycles to process $B$ micro-batches, resulting in a density of $B/(M+B-1)$. Asynchronous pipelining achieves a density of $1$, as all devices remain active during training.
Figure 3: Illustration of the dynamic of asynchronous pipeline. The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each device.
4 Convergence and Wall-clock Analysis
@@
This section provides the convergence guarantee of the proposed pipeline training algorithm on AIMC accelerators. We show that both synchronous and asynchronous pipelines enjoy the same convergence rate as the vanilla non-pipeline version. By increasing the computation density significantly, the pipeline parallelism provides notable speedup in terms of clock cycle complexity.
4.1 Assumptions on models, data, and AIMC accelerators
[…]
@@ 4.1 Assumptions on models, data, and AIMC accelerators
The sample $\xi_{k}$ are independently sampled from a distribution ${\mathcal{D}}$ over times $k\in[K]$, and has bounded variance, i.e., ${\mathbb{E}}_{\xi_{k}}[\nabla f(W_{k};\xi_{k})]=\nabla f(W_{k})$ and ${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}$.
@@
The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid and Tanh functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
@@
Note that the lower bounded loss implies the lower boundness of the objectives as well, i.e. $f(W;\xi)\geq f^{*}$, $\forall W\in{\mathbb{R}}^{D},x\in{\mathcal{X}}$. Aside from the assumptions, to avoid saturation during training, we also need proper scaling so that the saturation degree of the AIMC accelerators is bounded.
@@
There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$ norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max}<\tau,\forall k\in[K]$. The ratio $W_{\max}/\tau$ is referred to as saturation degree.
Assumption 5 requires the AIMC accelerators to have a sufficiently large active region to represent weights. Besides, on digital devices, the trainable weights always remain bounded in a small neighborhood of the initialization points during training [27, 28, 29], especially for overparameterized networks. Given $W_{\max}$ is sufficiently large to include this neighborhood, Assumption 5 holds. While it is possible to prove this assumption, we adopt it to make our results more concise and to focus on revealing the key insights on the effect of pipeline training.
Under the above assumptions, we prove the objective $f(W;\xi)$ is smooth.
@@
Under Assumption 25, the objective is $L_{f}$-smooth with respect to $W$, i.e. $\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$.
The proof of Lemma 1 is deferred to Appendix C. Lemma 1 enables us to adapt the tools from [30] for our analysis. However, due to the asymmetric bias of the AIMC accelerators, the convergence results for Analog SGD are still non-trivial.
4.2 Convergence rate of pipeline training algorithms
In this section, we provide the convergence rate of Analog SGD with pipelines. In non-convex optimization tasks like the DNN training, instead of measuring the objective (1), it is more common to measure the gradient square norm $\|\nabla f(W_{k})\|^{2}$.
Theorem 1 (Convergence rate, synchronous pipeline).

### v2

@@ II-C Synchronous and asynchronous pipelines
Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one accelerator and necessitates model parallelism. In model parallelism, the forward and backward computations are partitioned into a sequence of stages, which, in general, consist of a consecutive set of layers in the model. Each stage is mapped to one digital or analog accelerator. Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$. Without loss of generality, this paper assumes that one stage contains only one layer and the micro-batch size $B_{\text{micro}}=1$. Under this assumption, $\xi_{k,b}$ denotes the $b$-th micro-batch which contains only the $b$-th data. Consequently, the number of micro-batches is the mini-batch size, i.e., $B=B_{\text{mini}}$. This paper treats a stage as the minimal scheduling unit and investigates the time complexity across different scheduling strategies. To compare time costs, we define the basic time unit, clock cycle, as the minimum time required for all devices to process one micro-batch. Based on scheduling strategies, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining. Vanilla model parallelism without pipeline. Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data sample, as previously mentioned) for iteration $k$. Vanilla model parallelism processes micro-batches sequentially. In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$ according to (3), while in the backward pass stage $m$ achieves the backward signal $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage according to (4). The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stashed when the gradient for the $b$-th micro-batch is completed. After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method. Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}^{(m)}$, which iterates over $b\in[B_{\text{mini}}]$ by
@@
Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$. To fully release the potential of rank-update, AIMC accelerators follow (7) to update its weight, which in turn complicates the convergence analysis.
@@
Synchronous pipeline. Vanilla model parallelism is simple to implement. However, most devices remain idle, reducing system throughput. Existing work has already attempted to design an architecture that supports the pipeline training [30, 31]. They faithfully implement the iteration (7) but overlap the computing inside a mini-batch. Featuring synchronization at the end of a mini-batch, this stage schedule mechanism is called synchronous pipeline. In the vanilla non-pipeline version, stage $m$ remains idle after completing the forward computation of the $b$-th micro-batch until the backward signal returns from stage $m+1$. In the synchronous pipelining, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any. Similar to the vanilla model parallelism, the weights will be updated after processing all micro-batches. Consequently, for each mini-batch, most of the devices need to wait until the last micro-batch is processed. After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (7). To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch. During the pipeline filling and draining periods, parts of the devices are idle; see the middle of Fig. 2
Asynchronous pipeline. To fully utilize all devices, a more progressive strategy, asynchronous pipeline, is proposed. Instead of using mini-batch descent, asynchronous pipeline performs gradient descent on each micro-batch. Let $\xi_{k}$ denote the micro-batch for iteration $k$. In the forward pass, stage $m$ receives the output of the previous stage $x^{(m)}_{k}$ as input and outputs $x^{(m+1)}_{k}$, while in the backward pass, stage $m$ receives the backward signal $\delta^{(m)}_{k}$ from next stage and sends the error $\delta^{(m-1)}_{k}$ to its previous stage. Unlike the synchronous pipeline, the asynchronous pipeline updates the weights once $x^{(m)}_{k}$ and $\delta^{(m)}_{k}$ are ready. Without regular synchronization and filling/draining period, no device is idle during the training, maximizing the utilization. See Fig. 2 for the comparison of different parallel computing strategies.
@@
Despite its perfect resource usage, asynchronous pipeline parallelism suffers from the stale signal issue. Notice that in the asynchronous pipeline setting, the forward signal $x^{(m)}$ and backward signals $\delta^{(m)}$ are computed based on different stale weights. For example, after outputting forward signal $x^{(0)}$ for the $k$-th micro-batch, the first stage ($m=1$) receives the backward signal $\delta^{(0)}$ after 2(M1)2(M-1) clock cycles, wherein M1M-1 gradient updates happen. We introduce different symbols for precise description, which have tildes on top of them to indicate the staleness. Since there are MmM-m gradient updates between the forward and backward passes in stage $m$, the weight used for the forward pass has a delay of MmM-m. In contrast, the weights are updated by the backward pass, which means the weights used for the backward pass are the latest ones. The forward and backward passes of asynchronous pipeline training can be summarized as
Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$. (9)
Imperfect gradient-based training on AIMC accelerators. After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights. However, there is a bias due to the hardware imperfection [20]. Replacing $\Delta W_{k}$ with $\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ in (1), where $\alpha>0$ is a positive learning rate, we reach the update dynamics of weights
@@
$\textsf{Update}\quad W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
@@
In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
The main challenge coming from the AIMC training is the weight drift [20]. Consider the case where $W_{k}^{(m)}$ is already a non-zero critical point of $f(W)$ but not the one of $f(W;\xi)$, i.e. ${\mathbb{E}}_{\xi\sim{\mathcal{P}}}[\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)]=0$ but $\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)\neq 0$ for some $\xi$. After one iteration, the weight $W_{k}^{(m)}$ drifts because of the sample randomness
${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}\neq 0$ (11)
which prevents the weights from converging to the critical point, leading to an error.
III Convergence & Wall-clock Analysis
@@
This section shows that despite stale gradient issue and the update bias in AIMC accelerators, the asynchronous pipeline converges with the same sample complexity as the counterpart without pipeline or with synchronous pipeline. By increasing the computation density, asynchronous pipeline parallelism provides notable speedup in terms of wall-clock complexity.
III-A Assumptions
Before showing the complexity of the pipeline algorithms, we first introduce a series of assumptions on the properties of DNN structures, data distribution, as well as the AIMC accelerators. $\|\,\cdot\,\|$ is $\ell_{2}$-norm for vectors and Frobenious norm for matrices; $\|\,\cdot\,\|_{\infty}$ takes the element with maximal absolute value of vector or matrix.
@@ III-A Assumptions
The stochastic gradient is unbiased and has bounded variance $\sigma^{2}$.
@@
The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
@@
Aside from the assumptions, to avoid saturation during training, following the setting in [20], the saturation degree is assumed to be bounded.
@@
There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$. The ratio $W_{\max,\infty}/\tau$ is termed saturation degree.
From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range [20]. It is possible since the trainable weights always remain bounded in a small neighborhood of the initialization points during training [33, 34, 35], especially for overparameterized networks in digital training. Given $W_{\max,\infty}$ is sufficiently large to include this neighborhood, Assumption 5 holds. While it is possible to prove this assumption as in [36], we adopt it to focus on revealing key insights into the effect of pipeline training.
III-B Iteration complexity of pipeline training
Next, we present the convergence rate of analog SGD with pipelines. In non-convex optimization tasks like the DNN training, it is standard to use the stationary measure as $E_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]$. As a warm-up, we investigate the convergence of the synchronous pipeline, Analog-SGD-SP.
Theorem 1 (Iteration complexity, Analog-SGD-SP).
$E_{K}\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S$
where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}/\tau^{2}}{1-W_{\max,\infty}/\tau^{2}}$.
The proof of Theorem 1 is deferred to Appendix D. Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [37]. Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [20], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [20, 19, 22]. Next, we provide the iteration complexity of Analog-SGD-AP.
Theorem 2 (Iteration complexity, Analog-SGD-AP).
$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ (12)
The proof of Theorem 2 is deferred to Appendix E. Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which vanishes for a well-designed AIMC accelerator or can be mitigated by error-correction algorithms [20, 19]. Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence rate as the Analog-SGD-WOP [20].
Remark 1.
As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Theorem 3 in the Appendix F. Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [32].
III-C Proof sketch of Theorem 2
In this subsection, we highlight the key steps of the proof towards Theorem 2.
(Step I) Smoothness of the objective. Before analyzing the training dynamic, we show that the objective is smooth. Recall that $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$, we first show that under the assumptions of DNN structures, both $\delta^{(m)}$ and $x^{(m)}$ are smooth. Using slightly abusing notations, we denote the forward and backward signals computed by $W$ as $x^{(m)}$ and $\delta^{(m)}$, respectively; and denote the ones computed by $\tilde{W}$} as ${\tilde{x}}^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively. The following lemma show that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
Lemma 1 (Signal stability).
where $C_{x^{\prime}}^{(m)}$, $C_{z^{\prime}}^{(m)}$, and $C_{\delta^{\prime}}^{(m)}$ are constants that depend on the DNN architecture.
Based on this lemma, we prove the objective $f(W;\xi)$ is smooth.
@@ III-C Proof sketch of Theorem 2
Under Assumption 25, the objective is $L_{f}$-smooth with respect to $W$, i.e. $\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$, where $L_{f}$ is a constant depending on the DNN archetecture.
Lemma 2 enables us to adapt the tools from [37] to obtain the sufficient descent of the function value. However, the bias from the asymmetric update and stale signals from the asynchronous pipelining accumulates layer by layer within the DNN and eventually lead to an extra term in the upper bound. Next step we quantify the error of the stale signals.
(Step II) Error of stale signals. Similar to Step I, we begin from analyzing the error of $x^{(m)}$ and $\delta^{(m)}$. However, the stale signal involves the weights at different iterations and introduces additional errors. Intuitively, the error of forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal. It motivated us to bound the error by the delay. Since the error between latest $W^{(m^{\prime})}_{k}$ and the stale weight $W^{(m^{\prime})}_{k-(M-m)}$ can be decomposed as
$W^{(m^{\prime})}_{k}-W^{(m^{\prime})}_{k-(M-m)}=\!\!\!\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\!\!\!(W^{(m^{\prime})}_{k^{\prime}+1}-W^{(m^{\prime})}_{k^{\prime}})$, (16)
the following accumulated delay of layer $m$ can be used for analyzing the upper bound of the stale signals’ error
$\psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (17)
where $\Delta^{(m)}_{k}:=W^{(m)}_{k+1}-W^{(m)}_{k}$ is the difference between consecutive weights at iteration $k$.
Lemma 3 (Delay error of signals).
Under Assumption 24, the difference between the latest and stale signals generated by Analog-SGD-AP are bounded, i.e., $\forall m\in[M+1]$
$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\psi_{k},$ (18)
$\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\psi_{k},$ (19)
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\psi_{k},$ (20)
where $C_{x}^{(m)}$, $C_{z}^{(m)}$, and $C_{\delta}^{(m)}$ are constants that depends on the DNN architecture.
Remark 2.
The inequalities in Lemma 1 and 3 have a similar form, but they provide upper bounds for different error: in Lemma 1, two signals are computed on two arbitrary weights $W$ and $\tilde{W}$}, while in Lemma 3, two signals are generated by Analog-SGD-AP.
(Step III) Sufficient descent of the objective. Building on the results in Steps I and II, we will analyze the sufficient descent of the objective function. However, due to the asymmetric bias of the AIMC accelerators, the convergence results for analog SGD are still non-trivial.
Lemma 4 (Sufficient descent).
Under Assumption 15, the objective function $f(W_{k})$ satisfies